Write the chemical formula and answer the following questions below:



1. Barium nitride



2. Carbon tetrahydride



3. Iron (III) Chloride or Ferric chloride



4. Phosphorus pentoxide



5. Aluminium oxide



Questions:



1. What is the chemical formula of the compound?



2. Which among the compounds are ionic and covalent? How did you determine the difference?




1
Expert's answer
2021-11-08T08:27:40-0500

Part one


Compound Formula

Barium Nitride Ba3N2

Carbon tetrahydride CH4

Iron(iii) Chloride FeCl3

Phosphorus pentoxide P2O5

Aluminium Oxide Al2O3



Part two


Ionic bonds are formed when a complete transfer of electrons from one atom to another occurs

Usually its a bonding between a metal ( which loses valency electrons) and a non - metal which gain electrons


Covalent bonds are formed when two or more atoms combine which each atom donating equal number of electron(s) which are shared by the atoms involved.


In general, when two atoms have a small difference in their tendency to lose or gain electrons, sharing occurs leading to covalent bonding


"\therefore" The difference in negativity is another determinant on whether the bonding will be ionic or covalent.


If the electronegativity difference

("\Delta""EN)";

1) is less than 1.7 and non metals are involved, then the bonding is covalent


2) is greater than 1.7 and a metal is involved, the bonding is ionic


"\therefore" Barium Nitride, Iron (iii) Chloride and Aluminium Oxide forms ionic bondings.


The bondings involves a metal and a non-metal


"\Delta EN" ( Barium Nitride)= 3-0.9= 2.1


"\Delta EN" (Iron (iii) Chloride =

3.16-1.83= 1.33


"\Delta\>EN" (Aluminum Oxide)= 3.5-1.5 = 2


Carbon Tetrahydride and Phosphorus

Pentoxide have covalent bondings


Because the bondings involves two non metals


"\Delta" "EN" (Carbon Tetrahydride) = 2.55 - 2.2 = 0.35


"\Delta" "EN" (Phosphorus Pentoxide)= 3.5 - 2.19 = 1.24